# Theory of Irreversibility in Quantum Many-Body Systems

### v1

@@ Abstract
We address the longstanding challenge in quantum statistical mechanics of reconciling unitary dynamics with irreversible relaxation. In classical chaos, the unitary evolution operator develops Ruelle-Pollicott (RP) resonances inside the unit circle in the continuum limit, leading to mixing. In contrast, the quantum theory of many-body RP resonances and their link to irreversibility remain underdeveloped. Here, we relate the spectral form factor to the sum of autocorrelation functions and, in generic many-body lattice systems without conservation laws, argue that all quantum RP resonances converge inside the unit disk, highlighting the role of nonunitary and the thermodynamic limit. While we conjecture this picture to be general, we analytically prove the emergence of irreversibility in the random phase model (RPM), a paradigmatic Floquet quantum circuit model, in the limit of large local Hilbert space dimension. To this end, we couple it to local environments and compute the exact time evolution of autocorrelation functions, the dissipative form factor, and out-of-time-order correlation functions. Although valid for any dissipation strength, we then focus on weak dissipation to clarify the origin of irreversibility in unitary systems. When the dissipationless limit is taken after the thermodynamic limit, the unitary quantum map develops an infinite tower of RP resonances—chaotic systems display so-called anomalous relaxation. We identify the exact RP resonances in the RPM and prove that the same RP resonances are obtained from operator truncation. We also show that the OTOC in the RPM can undergo a two-stage relaxation and that during the second stage, the approach to the stationary value is again controlled by the leading RP resonance. Finally, we demonstrate how conservation laws, many-body localization, and nonlocal interactions merge the leading RP resonance into the unit circle, thereby suppressing anomalous relaxation. Our findings clarify the microscopic origin of macroscopic relaxation in quantum systems and emphasize the role of RP resonances as a largely unexplored signature of many-body quantum chaos.
@@ I Introduction
The same mechanism is at play in quantum systems [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. The density matrix of an ensemble of quantum systems evolves under a quantum channel, which is unitary for closed systems. Yet, correlation functions decay, and, correspondingly, some quantum analog of RP resonances must be hidden in the spectrum of the quantum channel. A first hurdle to this approach is that the spectrum of a quantum channel is not dense even if the system is chaotic—a finite-size quantum chaotic system has a discrete spectrum with eigenvalues correlated according to random matrix theory (RMT) [18, 19]. Accordingly, to obtain a dense spectrum one needs to take the thermodynamic limit. Given the practical impossibility of taking the strict thermodynamic limit, alternative mechanisms have been devised that provide a “seed” for quantum RP resonances, making them visible in large but finite systems. Prosen [9, 10, 11] first proposed coarse-graining in operator space, which renders the quantum channel nonunitary. With an appropriately designed algorithm [9, 14], making the coarse-graining finer amounts to simultaneously taking the thermodynamic and unitary limits in the correct order and one can observe a quick convergence of eigenvalues to points inside the unit disk—these are the quantum RP resonances. Remarkably, this procedure provided not only fundamental insight into the structure of many-body quantum systems but also a powerful numerical tool, as the decay of correlation functions converged faster to the thermodynamic limit with the coarse-graining procedure. The same idea is also behind subsequent efficient numerical methods [20, 21].
@@
More recently, it was realized that explicitly coupling the system of interest to an external bath provides an alternative seed for quantum many-body RP resonances, with the advantage of also elucidating the physics of the weak-dissipation limit of chaotic open many-body quantum systems [22, 23, 24, 13, 25]. These systems retain a finite (order one) relaxation rate even when the dissipation strength vanishes, provided that the thermodynamic limit is taken before the weak-dissipation limit—a phenomenon dubbed anomalous relaxation in Ref. [26]. Anomalous relaxation was first observed in the dissipative Sachdev-Ye-Kitaev model [22, 26] and, later, in a dissipative quantum spin-liquid model [24], the driven-dissipative transverse-field Ising model [13], and an open quantum circuit [25]. It was conjectured to be a generic feature of chaotic systems and absent in integrable ones [26], with Mori [13] formalizing these ideas and connecting them to quantum many-body RP resonances. We note that Ref. [26] also put forward a possible holographic interpretation of anomalous relaxation—and, therefore, quantum many-body RP resonances—in gravity, see also Ref. [27].
@@
Despite all this recent progress, the theory of quantum many-body RP resonances is still rudimentary, especially when compared with its established classical counterpart. The existence of quantum RP resonances can be rigorously established in certain single-particle [28, 12] and quadratic many-body [8] systems but most of the scarce evidence in interacting many-body systems is heuristic or numerical. The notable exception is the dissipative random phase model (DRPM), a quantum-chaotic dissipative quantum circuit model introduced by the same authors of the present paper in Ref. [25]. In the limit of large local Hilbert-space dimension, this model is exactly solvable for arbitrary system size and dissipation strength and constitutes, therefore, a perfect testbed to further develop the theory of irreversibility of many-body quantum systems. In this paper, we proceed in this direction with four distinct goals in mind.
@@
Second, we exploit the exact solvability of the DRPM in the limit of large local Hilbert space dimension to perform the first analytical computation of a full quantum many-body RP resonance spectrum. We obtain not only the leading RP resonance [25] but an infinite tower of RP resonances that rule the decay of all possible autocorrelation functions in the closed-system limit of the DRPM, the random phase model (RPM) [29]. Furthermore, the analytical control we have over the model also provides a clear physical mechanism for the emergence of the quantum many-body RP resonances, namely, the same domain walls responsible for the Thouless physics in the RPM [29]—the physics before the onset of universal random-matrix behavior. We conjecture this mechanism to be universal in quantum many-body systems.
Third, we study a so-far unexplored dynamical regime of many-body quantum chaos, the Ruelle regime. Arguably, the two most striking features of chaos are sensitivity to initial conditions and mixing. If a quantum system has a well-defined semiclassical regime, the former manifests as an early-time exponential growth in out-of-time-order correlators (OTOCs), which defines a quantum Lyapunov exponent. The Lyapunov regime has been extensively investigated for both single-particle [30, 31, 32, 12] and many-body systems [33, 34], but many systems without an effective semiclassical parameter completely lack this behavior. On the contrary, the decay of correlation functions is a universal feature of all chaotic quantum systems—the quantum version of mixing. To illustrate this point, we analytically demonstrate that, while the OTOC of the RPM has no Lyapunov regime, the same quantum RP resonances as before rule the saturation of the OTOC (and its decay in the presence of dissipation), which is identified as the Ruelle regime of the model.
@@ II.2 Quantum Ruelle-Pollicott resonances
Compared to the classical RP resonances, much less has been understood about their quantum counterparts, especially in many-body settings. The first attempt to generalize the classical RP resonances to many-body quantum settings was made by Prosen in Ref. [9] where a course-graining procedure using operator truncation was introduced to define a nonunitary time-evolution operator. This procedure was inspired by analytic continuation in the classical case, and the eigenvalues of the nonunitary operator were then identified as the quantum RP resonances, which dictate the late-time dynamics of the system. Recently, another way of defining the quantum RP resonances, this time paralleling the stochastization of the Liouville operator, was proposed by Mori [13]. Namely, it was suggested that the eigenvalues of the time-evolution operator in the presence of weak dissipation are related to quantum RP resonances and, in particular, they coincide precisely with the RP resonances in the limit of vanishing dissipation after taking the thermodynamic limit. In this paper, we shall primarily use the second approach to define the quantum RP resonances. We shall demonstrate that the above two approaches follow from the same physical mechanism and, hence, give rise to the same RP resonances.
@@ III Case study: the (dissipative) random phase model
In this section, we analytically compute, for arbitrary onsite quantum channels, different dynamical quantities that characterize the evolution of the DRPM, defined in Sec. II.4.1: autocorrelation functions of local operators, the dissipative form factor, the truncated form factor, and out-of-time-order correlators. We are particularly interested in the late-time decay of these quantities and extracting quantum many-body RP resonances from them.
@@ III.4 OTOCs and their relaxation
So far we have discussed how the quantum RP resonances govern the decay of the ensemble-averaged correlation functions. Now we turn to the OTOCs, which are another important probe of quantum chaos. We demonstrate that quantum RP resonances also control their saturation to the stationary value one using the exact OTOCs of the large-$q$ RPM obtained in Ref. [65]. We then show that dissipation induces an overall decay of the OTOCs in the DRPM, which is again dictated by the leading RP resonance.
@@ V Discussion
In this paper, we laid down a general theory that describes how the quantum many-body analogs of RP resonances govern the emergence of irreversible dynamics in isolated quantum many-body systems. Building upon a recent work by the same authors of the present paper [25], we first explained why it is possible to extract the quantum RP resonances from the long-time behavior of the DFF by identifying it with the sum of autocorrelation functions whose asymptotic decay in time is controlled by the RP resonances on the $O(L)$ timescale, after which a linear ramp emerges [65]. In particular, we clarified the importance of taking the thermodynamic limit before the dissipationless limit in obtaining the leading RP resonance from the DFF; this, in turn, gives rise to a nonzero Liouvillian gap of the system in the limit of vanishing dissipation [13], a phenomenon dubbed anomalous relaxation [26]. At the level of the sum of autocorrelation functions, dissipation suppresses the ramp of the SFF, which starts at shorter timescales for highly nonlocal operators. This observation also allowed us to demonstrate that operator truncation, which is another approach to defining the quantum RP resonances [9, 10, 11, 14, 15], is essentially equivalent to the above prescription using weak dissipation and therefore yields the same set of quantum RP resonances. To establish these claims in a concrete setting, we fully worked out the random phase model [29] and the dissipative random phase model [25] at large $q$.
To further substantiate our claim that the quantum RP resonances dictate the relaxation of quantum many-body dynamics, we looked into the OTOC for the RPM at large $q$ and showed that with PBCs it undergoes two-stage relaxation. In the second stage, the relaxation is fully controlled by the leading RP resonance, as expected. Furthermore, we evaluated the OTOC in the large-$q$ DRPM exactly and demonstrated that it again decays with the leading RP resonance in the dissipationless limit after taking the thermodynamic limit, similarly to the DFF. Finally, we examined several scenarios in which quantum RP resonances are constrained to be on the unit circle, including the case where the system has a $U(1)$ charge, both in solvable models and numerically.
@@
Another important direction that warrants further study is to look into the system with conserved charges, in particular Hamiltonian systems, in more detail. While quantum circuits with a $U(1)$-charge have been routinely used as toy models that mimic Hamiltonian systems in certain aspects, it remains unclear whether the quantum RP resonances also behave in similar ways in both systems. Indeed, in Ref. [13], it was discovered that the RP resonance in the kicked Ising model, which conserves the energy, is proportional to $L^{-1}$ despite the naive expectation that it decays diffusively, i.e., $L^{-2}$. A systematic comparison between systems that conserve a $U(1)$ charge and energy is, therefore, highly desired.
VI Acknowledgements

### v2

@@ Abstract
We address the longstanding challenge in quantum many-body theory of reconciling unitary dynamics with irreversible relaxation. In classical chaos, the unitary evolution operator develops Ruelle-Pollicott (RP) resonances inside the unit circle in the continuum limit, leading to mixing. In the semiclassical limit, chaotic single-particle quantum systems relax with the same RP resonances. In contrast, the theory of quantum many-body RP resonances and their link to irreversibility remain underdeveloped. Here, we relate the spectral form factor to the sum of autocorrelation functions and, in generic many-body lattice systems without conservation laws, argue that all quantum many-body RP resonances converge inside the unit disk, highlighting the role of nonunitary and the thermodynamic limit. While we conjecture this picture to be general, we analytically prove the emergence of irreversibility in the random phase model (RPM), a paradigmatic Floquet quantum circuit model, in the limit of large local Hilbert space dimension. To this end, we couple it to local environments and compute the exact time evolution of autocorrelation functions, the dissipative form factor, and out-of-time-order correlation functions (OTOCs). Although valid for any dissipation strength, we then focus on weak dissipation to clarify the origin of irreversibility in unitary systems. When the dissipationless limit is taken after the thermodynamic limit, the unitary quantum map develops an infinite tower of decaying RP resonances—chaotic systems display so-called anomalous relaxation. We identify the exact RP resonances in the RPM and prove that the same RP resonances are obtained from operator truncation. We also show that the OTOC in the RPM can undergo a two-stage relaxation and that during the second stage, the approach to the stationary value is again controlled by the leading RP resonance. Finally, we demonstrate how conservation laws, many-body localization, and nonlocal interactions merge the leading RP resonance into the unit circle, thereby suppressing anomalous relaxation. Our findings clarify the microscopic origin of macroscopic relaxation in quantum systems and emphasize the role of RP resonances as a largely unexplored signature of many-body quantum chaos.
@@ I Introduction
The same mechanism is at play in quantum systems [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. The density matrix of an ensemble of quantum systems evolves under a quantum channel, which is unitary for closed systems. Yet, correlation functions decay, and, correspondingly, some quantum analog of RP resonances must be hidden in the spectrum of the quantum channel. A first hurdle to this approach is that the spectrum of a quantum channel is not dense even if the system is chaotic—a finite-size quantum chaotic system has a discrete spectrum with eigenvalues correlated according to random matrix theory (RMT) [22, 23]. Accordingly, to obtain a dense spectrum, one needs to take either the semiclassical [8, 9, 10, 11, 12] or the thermodynamic limit [13, 14, 15, 16, 17, 18, 19, 20, 21]. In single-particle system, the former recovers the classical RP resonances; in many-body systems, the latter is required. Given the practical impossibility of taking the strict thermodynamic limit, alternative mechanisms have been devised that provide a “seed” for quantum RP resonances, making them visible in large but finite systems. Prosen [13, 14, 15] first proposed coarse-graining in operator space, which renders the quantum channel nonunitary. With an appropriately designed algorithm [13, 18], making the coarse-graining finer amounts to simultaneously taking the thermodynamic and unitary limits in the correct order and one can observe a quick convergence of eigenvalues to points inside the unit disk—these are the quantum RP resonances. Remarkably, this procedure provided not only fundamental insight into the structure of many-body quantum systems but also a powerful numerical tool, as the decay of correlation functions converged faster to the thermodynamic limit with the coarse-graining procedure. The same idea is also behind subsequent efficient numerical methods [24, 25].
@@
More recently, it was realized that explicitly coupling the system of interest to an external bath provides an alternative seed for quantum RP resonances, with the advantage of also elucidating the physics of the weak-dissipation limit of chaotic open many-body quantum systems [26, 27, 28, 16, 17]. These systems retain a finite (order one) relaxation rate even when the dissipation strength vanishes, provided that the thermodynamic limit is taken before the weak-dissipation limit—a phenomenon dubbed anomalous relaxation in Ref. [29]. Anomalous relaxation was first observed in the dissipative Sachdev-Ye-Kitaev model [26, 29] and, later, in a dissipative quantum spin-liquid model [28], the driven-dissipative transverse-field Ising model [16], and an open quantum circuit [17]. It was conjectured to be a generic feature of chaotic systems and absent in integrable ones [29], with Mori [16] formalizing these ideas and connecting them to quantum RP resonances. We note that Ref. [29] also put forward a possible holographic interpretation of anomalous relaxation—and, therefore, quantum RP resonances—in gravity, see also Ref. [30].
@@
Despite all this recent progress, the theory of quantum many-body RP resonances is still rudimentary. The existence of quantum RP resonances can be rigorously established in certain single-particle systems [31, 9, 10, 11, 12], but most of the scarce evidence in interacting many-body systems is heuristic or numerical. The notable exception is the dissipative random phase model (DRPM), a quantum-chaotic dissipative quantum circuit model introduced by the same authors of the present paper in Ref. [17]. In the limit of large local Hilbert space dimension, this model is exactly solvable for arbitrary system size and dissipation strength and constitutes, therefore, a perfect testbed to further develop the theory of irreversibility of many-body quantum systems. In this paper, we proceed in this direction with four distinct goals in mind.
@@
Second, we exploit the exact solvability of the DRPM in the limit of large local Hilbert space dimension to perform the first analytical computation of a full quantum many-body RP resonance spectrum. We obtain not only the leading RP resonance [17] but an infinite tower of RP resonances that rule the decay of all possible autocorrelation functions in the closed-system limit of the DRPM, the random phase model (RPM) [32]. Furthermore, the analytical control we have over the model also provides a clear physical mechanism for the emergence of the quantum RP resonances, namely, the same domain walls responsible for the Thouless physics in the RPM [32]—the physics before the onset of universal random-matrix behavior. We conjecture this mechanism to be universal in quantum many-body systems.
Third, we study a so-far unexplored dynamical regime of many-body quantum chaos, the Ruelle regime. Arguably, the two most striking features of chaos are sensitivity to initial conditions and mixing. If a quantum system has a well-defined semiclassical regime, the former manifests as an early-time exponential growth in out-of-time-order correlators (OTOCs), which defines a quantum Lyapunov exponent. The Lyapunov regime has been extensively investigated for both single-particle [33, 34, 35, 12] and many-body systems [36, 37], but many systems without an effective semiclassical parameter completely lack this behavior. On the contrary, the decay of correlation functions is a universal feature of all chaotic quantum systems—the quantum version of mixing. To illustrate this point, we analytically demonstrate that, while the OTOC of the RPM has no Lyapunov regime, the same quantum RP resonances as before rule the saturation of the OTOC [38] (and its decay in the presence of dissipation), which is identified as the Ruelle regime of the model.
@@ II.2 Quantum Ruelle-Pollicott resonances
The previous construction holds, essentially unaltered, for single-particle quantum systems, e.g., in low-dimensional quantum maps [9, 10, 11, 12] or quantum billiards [42, 43]. In the semiclassical limit, correlations decay with a rate given by the RP resonances of the limiting classical dynamical system. In contrast, truly quantum RP resonances occur only for quantum systems in the thermodynamic limit, even without any classical limit; how these emerge from the unitary dynamics of isolated systems is the focus of the rest of the paper.
The first attempt to generalize the concept of RP resonances to many-body quantum settings was made by Prosen in Ref. [13] where a course-graining procedure using operator truncation was introduced to define a nonunitary time-evolution operator. This procedure was inspired by analytic continuation in the classical case, and the eigenvalues of the nonunitary operator were then identified as the quantum RP resonances, which dictate the late-time dynamics of the system. Recently, another way of defining the quantum RP resonances, this time paralleling the stochastization of the Liouville operator, was proposed by Mori [16]. Namely, it was suggested that the eigenvalues of the time-evolution operator in the presence of weak dissipation are related to quantum RP resonances and, in particular, they coincide precisely with the RP resonances in the limit of vanishing dissipation after taking the thermodynamic limit. In this paper, we shall primarily use the second approach to define the quantum RP resonances. We shall demonstrate that the above two approaches follow from the same physical mechanism and, hence, give rise to the same RP resonances.
@@ III Case study: the (dissipative) random phase model
In this section, we analytically compute, for arbitrary onsite quantum channels, different dynamical quantities that characterize the evolution of the DRPM, defined in Sec. II.4.1: autocorrelation functions of local operators, the dissipative form factor, the truncated form factor, and out-of-time-order correlators. We are particularly interested in the late-time decay of these quantities and extracting quantum RP resonances from them.
@@ III.4 OTOCs and their relaxation
So far we have discussed how the quantum RP resonances govern the decay of the ensemble-averaged autocorrelation functions. Now we turn to the OTOCs, which are another important probe of quantum chaos. We demonstrate that quantum RP resonances also control their saturation to the stationary value one using the exact OTOCs of the large-$q$ RPM obtained in Ref. [71]. We then show that dissipation induces an overall decay of the OTOCs in the DRPM, which is again dictated by the leading RP resonance.
@@ V Discussion
In this paper, we laid down a general theory that describes how the quantum many-body analogs of RP resonances govern the emergence of irreversible dynamics in isolated quantum many-body systems. Building upon a recent work by the same authors of the present paper [17], we first explained why it is possible to extract the quantum RP resonances from the long-time behavior of the DFF by identifying it with the sum of autocorrelation functions whose asymptotic decay in time is controlled by the RP resonances on the $O(L)$ timescale, after which a linear ramp emerges [71]. In particular, we clarified the importance of taking the thermodynamic limit before the dissipationless limit in obtaining the leading RP resonance from the DFF; this, in turn, gives rise to a nonzero Liouvillian gap of the system in the limit of vanishing dissipation [16], a phenomenon dubbed anomalous relaxation [29]. At the level of the sum of autocorrelation functions, dissipation suppresses the ramp of the SFF, which starts at shorter timescales for highly nonlocal operators. This observation also allowed us to demonstrate that operator truncation, which is another approach to defining quantum RP resonances [13, 14, 15, 18, 19], is essentially equivalent to the above prescription using weak dissipation and therefore yields the same set of quantum RP resonances. To establish these claims in a concrete setting, we fully worked out the random phase model [32] and the dissipative random phase model [17] at large $q$.
To further substantiate our claim that the RP resonances dictate the relaxation of quantum many-body dynamics, we looked into the OTOC for the RPM at large $q$ and showed that with PBCs it undergoes two-stage relaxation. In the second stage, the relaxation is fully controlled by the leading RP resonance, as expected. Furthermore, we evaluated the OTOC in the large-$q$ DRPM exactly and demonstrated that it again decays with the leading RP resonance in the dissipationless limit after taking the thermodynamic limit, similarly to the DFF. Finally, we examined several scenarios in which quantum RP resonances are constrained to be on the unit circle, including the case where the system has a $U(1)$ charge, both in solvable models and numerically.
@@
Another important direction that warrants further study is to look into the system with conserved charges, in particular Hamiltonian systems, in more detail. While quantum circuits with a $U(1)$-charge have been routinely used as toy models that mimic Hamiltonian systems in certain aspects, it remains unclear whether the quantum RP resonances behave similarly in both systems. Indeed, in Ref. [16], it was discovered that the RP resonance in the kicked Ising model, which conserves the energy, is proportional to $L^{-1}$ despite the naive expectation that it decays diffusively, i.e., $L^{-2}$. A systematic comparison between systems that conserve a $U(1)$ charge and energy is, therefore, highly desired.
Acknowledgements